OpenStudy (anonymous):

If 10.1 of Fe are added to a solution of Copper (Cu) sulfate, how many grams of solid copper would form? 2Fe + 3CuSO4 -> FFe2(SO4)3 + 3Cu

OpenStudy (anonymous):

@chemIII

OpenStudy (anonymous):

@the5the

OpenStudy (anonymous):

\[ \Big( \frac{10.1 \ g \ \text{Fe}}{} \Big) \Big( \frac{1 \ \text{mol} \ \text{Fe}}{55.8 \ g \ \text{Fe}} \Big)\Big( \frac{3 \ \text{mol} \ \text{Cu} }{2 \ \text{mol} \ \text{Fe}} \Big)\Big( \frac{63.5 \ g \ \text{Cu}}{1 \ \text{mol} \ \text{Cu}} \Big)=\text{grams solid copper}\]